# Partial information decomposition as information bottleneck

### v1

@@ V Discussion
We show that for any $Q$ that achieves $I(Q;S|Y)\leq R$, there is another $Q^{\prime}$ with cardinality $\left|\mathcal{Q}^{\prime}\right|\leq\sum_{s}\left|\mathcal{X}_{s}\right|+1$ that satisfies $I(Q^{\prime};S|Y)\leq R$ and $I(Q^{\prime};Y|S)\geq I(Q;Y|S)$.
Consider any joint distribution $p_{QSZY}$ from (10) that achieves $I(Q;S|Y)\leq R$. Fix the corresponding conditional distribution $p_{SZ|Q}$, and note that it also determines the conditional distributions
$p_{YSZ|Q}(y,s,x|q)$ $=p_{Y|SZ}(y|s,z)p_{SZ|Q}(s,z|q)$
$=\frac{\nu_{S}(s)p_{X_{s}|Y}(z|y)p_{Y}(y)}{p_{SZ}(s,z)}p_{SZ|Q}(s,z|q)$
$p_{Y|SQ}(y|s,q)$ $=\frac{\sum_{z}p_{YSZ|Q}(y,s,z|q)}{\sum_{z,y^{\prime}}p_{YSZ|Q}(y^{\prime},s,z|q)}$
$p_{S|YQ}(s|y,q)$ $=\frac{\sum_{z}p_{YSZ|Q}(y,s,z|q)}{\sum_{z,s^{\prime}}p_{YSZ|Q}(y,s^{\prime},z|q)}$
Next, consider the following linear program,
@@
$\sum_{q}\omega_{Q^{\prime}}(q)H(p_{S|Y,Q=q})=H(p_{S|YQ})$ (32)
where $\Delta$ is the $|\mathcal{Q}|$-dimensional unit simplex, $D$ is KL divergence, and $H(\cdot)$ is Shannon entropy. We also used the notation $H(p_{S|Y,Q=q})=-\sum_{y,s}p_{YS|Q}(y,s|q)\ln p_{S|YQ}(s|y,q)$. The first set of constraints (31) guarantees that $\omega_{Q^{\prime}}p_{YSZ|Q}$ belongs to the family (10), and in particular that marginal over $(S,Z,Y)$ is $\nu_{S}(s)p_{X_{s}|Y}(x|y)p_{Y}(y)$. There are $\sum_{s}\left|\mathcal{X}_{s}\right|$ possible outcomes of $(s,z)$, but $\sum_{s,z}p_{SZ}(s,z)=1$ by conservation of probability, so effectively this imposes $\sum_{s}\left|\mathcal{X}_{s}\right|-1$ constraints. The last constraint (32) guarantees that $H(p_{S|YQ^{\prime}})=H(p_{S|YQ})$, hence
$I(Q^{\prime};S|Y)$ $=H(p_{S|Y})-H(p_{S|YQ^{\prime}})$
$=H(p_{S|Y})-H(p_{S|YQ})=I(Q;S|Y)\leq R.$
Eq. (30) involves a maximization of a linear function over the simplex subject to $\sum_{s}\left|\mathcal{X}_{s}\right|$ hyperplane constraints. The feasible set is compact, and the maximum will be achieved at one of the extreme points of the feasible set. By Dubin’s Theorem [57], any extreme point of this feasible set can be expressed as a convex combination of at most $\sum_{s}\left|\mathcal{X}_{s}\right|+1$ extreme points of $\Delta$. Thus, the maximum can be achieved by a marginal distribution $\omega_{Q^{\prime}}$ with support on at most $\sum_{s}\left|\mathcal{X}_{s}\right|+1$ outcomes. This distribution satisfies
$\sum_{q}\omega_{Q^{\prime}}(q)D(P_{Y|SQ=q}\|P_{Y|S})\\ \geq\sum_{q}p_{Q}(q)D(P_{Y|SQ=q}\|P_{Y|S})$
since the actual marginal $p_{Q}$ is an element of the feasible set. Finally, note that
@@
For a finite-dimensional system, we may restrict the optimization problem in Theorem 1 to $Q$ with cardinality $|\mathcal{Q}|\leq\sum_{s}\left|\mathcal{X}_{s}\right|+1$ (Theorem 3). In this case, the feasible set can be restricted to a compact set and the objective is continuous, therefore the supremum will be achieved.
Now consider a tuple of random variables $(S,Z,Y,Q)$ that obey the Markov conditions $S-Y-Z$ and $Q-(S,Z)-Y$. Suppose that $Q$ achieves the supremum in Theorem 1 for a given $R>0$,
@@
Consider also sequence of random variables $(S_{k},Z_{k},Y_{k},Q_{k})$ for $k=1,2,3\dots$. Suppose that each tuple has the same outcomes as $(S,Z,Y,Q)$, and it obeys the Markov conditions $S_{k}-Y_{k}-Z_{k}$ and $Q_{k}-(S_{k},Z_{k})-Y_{k}$. Let $I_{\mathrm{\text{RB}}}^{k}(R)$ indicate the redundancy bottleneck defined in Theorem 1 for random variables $Z_{k},Y_{k},S_{k}$, and suppose that $Q_{k}$ achieves the optimum for problem $k$,
@@
Given our assumption that $p_{SZ}$ has full support, we can always take $k$ sufficiently large so that each $p_{S_{k}Z_{k}}$ has full support.
@@
We define the random variable $Q_{k}^{\prime}$ that obeys the Markov condition $Q_{k}^{\prime}-(S,Z)-Y$, with conditional distribution
@@
This conditional distribution is always well-defined, given that $p_{S_{k}Z_{k}}$ has the same support as $p_{SZ}$. By assumption, $p_{S_{k}Z_{k}Y_{k}}\to p_{SZY}$, therefore
@@
Conditional mutual information is uniformly continuous due to uniform continuity of entropy [Theorem 17.3.3, 58]. Therefore
00 0 $=\lim_{k}I(Q_{k};S_{k}|Y_{k})-I(Q_{k}^{\prime};S|Y)$
$\leq R-\lim_{k}I(Q_{k}^{\prime};S|Y),$ (37)
where we used Eq. (34). We also define another random variable $Q_{k}^{\alpha}$ which also obeys the Markov condition $Q_{k}^{\alpha}-(S,Z)-Y$, whose conditional distribution is defined in terms of the convex mixture
@@
Here $p_{U}(q)=1/|\mathcal{Q}|$ is a uniform distribution over an auxiliary independent random variable $U$ with outcomes $\mathcal{Q}$. Using standard techniques [59], it can be shown that the conditional mutual information $I(Q_{k}^{\alpha};S|Y)$ is convex,
@@
Combining Eqs. (37), (38), and $R>0$ implies that $\alpha_{k}\to 1$, so $p_{Q_{k}^{\alpha}SZY}\to p_{Q_{k}^{\prime}SZY}$. Combining with Eqs. (39), (36), (34), along with continuity of conditional mutual information, gives
$I_{\text{RB}}(R)$ $\geq\lim_{k}I(Q_{k}^{\alpha};Y|S)$
$=\lim_{k}I(Q_{k}^{\prime};Y|S)$
$=\lim_{k}I(Q_{k};Y_{k}|S_{k})=\lim_{k}I_{\cap}^{k}(R).$
@@
00 0 $=\lim_{k}I(Q;S|Y)-I(Q_{k}^{\prime\prime};S_{k}|Y_{k})$
$\leq R-\lim_{k}I(Q_{k}^{\prime\prime};S_{k}|Y_{k}).$ (42)
@@
Using convexity of conditional mutual information, $I(U;S_{k};Y_{k})=0$ and the definition of $\alpha_{k}^{\prime}$, we have
@@
Combining Eqs. (42), (43), and $R>0$ implies $\alpha_{k}^{\prime}\to 1$, therefore $p_{Q_{k}^{\alpha\prime}S_{k}X_{k}Y_{k}}-p_{Q_{k}^{\prime\prime}S_{k}X_{k}Y_{k}}\to 0$. Combining with Eqs. 44, (41), (33), along with continuity of conditional mutual information, gives
$\lim_{k}I_{\cap}^{k}(R)\geq$ $\lim_{k}I(Q_{k}^{\alpha\prime};Y_{k}|S_{k})$
$=$ $\lim_{k}I(Q_{k}^{\prime\prime};Y_{k}|S_{k})$
$=$ $I(Q;Y|S)=I_{\text{RB}}(R).$
Acknowledgements.
I thank Nihat Ay, Daniel Polani, Fernando Rosas, and especially André Gomes for useful feedback. I also thank the organizers of the “Decomposing Multivariate Information in Complex Systems” (DeMICS 23) workshop at the Max Planck Institute for the Physics of Complex Systems (Dresden, Germany), which inspired this work.

### v2

@@ Acknowledgements.
We show that, for any $Q$ that achieves $I(Q;S|Y)\leq R$, there is another $Q^{\prime}$ with cardinality $\left|\mathcal{Q}^{\prime}\right|\leq\sum_{s}\left|\mathcal{X}_{s}\right|+1$ that satisfies $I(Q^{\prime};S|Y)\leq R$ and $I(Q^{\prime};Y|S)\geq I(Q;Y|S)$.
Consider any joint distribution $p_{QSZY}$ from Eq. (10) that achieves $I(Q;S|Y)\leq R$, and let $\mathcal{Q}$ be the corresponding set of outcomes of $Q$. Fix the corresponding conditional distribution $p_{SZ|Q}$, and note that it also determines the conditional distributions:
$p_{YSZ|Q}(y,s,z|q)$ $=p_{Y|SZ}(y|s,z)p_{SZ|Q}(s,z|q)$ (35)
$=\frac{\nu_{S}(s)p_{X_{s}|Y}(z|y)p_{Y}(y)}{p_{SZ}(s,z)}p_{SZ|Q}(s,z|q)$ (36)
$p_{Y|SQ}(y|s,q)$ $=\frac{\sum_{z}p_{YSZ|Q}(y,s,z|q)}{\sum_{z,y^{\prime}}p_{YSZ|Q}(y^{\prime},s,z|q)}$ (37)
$p_{S|YQ}(s|y,q)$ $=\frac{\sum_{z}p_{YSZ|Q}(y,s,z|q)}{\sum_{z,s^{\prime}}p_{YSZ|Q}(y,s^{\prime},z|q)}$ (38)
Next, consider the following linear program:
@@
$\sum_{q}\omega_{Q^{\prime}}(q)H({S|Y,Q=q})=H({S|Y,Q})$ (41)
where $\Delta$ is the $|\mathcal{Q}|$-dimensional unit simplex, and we use the notation $H({S|Y,Q=q})=-\sum_{y,s}p_{YS|Q}(y,s|q)\ln p_{S|YQ}(s|y,q)$. The first set of constraints (40) guarantees that $\omega_{Q^{\prime}}p_{YSZ|Q}$ belongs to the family (10) and, in particular, that the marginal over $(S,Z,Y)$ is $\nu_{S}(s)p_{X_{s}|Y}(z|y)p_{Y}(y)$ (see Eq. (36)). There are $\sum_{s}\left|\mathcal{X}_{s}\right|$ possible outcomes of $(s,z)$, but $\sum_{s,z}p_{SZ}(s,z)=1$ by the conservation of probability. Therefore, Eq. (40) effectively imposes $\sum_{s}\left|\mathcal{X}_{s}\right|-1$ constraints. The last constraint (41) guarantees that $H({S|Y,Q^{\prime}})=H({S|Y,Q})$; hence,
$I(Q^{\prime};S|Y)$ $=H({S|Y})-H({S|Y,Q^{\prime}})$
$=H({S|Y})-H({S|Y,Q})=I(Q;S|Y)\leq R.$
Eq. (39) involves a maximization of a linear function over the simplex, subject to $\sum_{s}\left|\mathcal{X}_{s}\right|$ hyperplane constraints. The feasible set is compact, and the maximum is achieved at one of the extreme points of the feasible set. By Dubin’s theorem [56], any extreme point of this feasible set can be expressed as a convex combination of at most $\sum_{s}\left|\mathcal{X}_{s}\right|+1$ extreme points of $\Delta$. Thus, the maximum is achieved by a marginal distribution $\omega_{Q^{\prime}}$ with support on at most $\sum_{s}\left|\mathcal{X}_{s}\right|+1$ outcomes. This distribution satisfies:
$\sum_{q}\omega_{Q^{\prime}}(q)D(P_{Y|SQ=q}\|P_{Y|S})\geq\sum_{q}p_{Q}(q)D(P_{Y|SQ=q}\|P_{Y|S})$
since the actual marginal distribution $p_{Q}$ is an element of the feasible set. Finally, note that
@@
For a finite-dimensional system, we may restrict the optimization problem in Theorem 1 to $Q$ with cardinality $|\mathcal{Q}|\leq\sum_{s}\left|\mathcal{X}_{s}\right|+1$ (Theorem 3). In this case, the feasible set can be restricted to a compact set, and the objective is continuous; therefore, the maximum will be achieved.
Now, consider a tuple of random variables $(S,Z,Y,Q)$ that obey the Markov conditions $S-Y-Z$ and $Q-(S,Z)-Y$. Suppose that $Q$ achieves the maximum in Theorem 1 for a given $R>0$:
@@
Consider also a sequence of random variables $(S_{k},Z_{k},Y_{k},Q_{k})$ for $k=1,2,3\dots$ where each tuple has the same outcomes as $(S,Z,Y,Q)$ and obeys the Markov conditions $S_{k}-Y_{k}-Z_{k}$ and $Q_{k}-(S_{k},Z_{k})-Y_{k}$. Let $I_{\mathrm{\text{RB}}}^{k}(R)$ indicate the redundancy bottleneck defined in Theorem 1 for random variables $Z_{k},Y_{k},S_{k}$, and suppose that $Q_{k}$ achieves the optimum for problem $k$:
@@
First, observe that given our assumption that $p_{SZ}$ has full support, we can always take $k$ sufficiently large so that each $p_{S_{k}Z_{k}}$ has full support. Next, we define the random variable $Q_{k}^{\prime}$ that obeys the Markov condition $Q_{k}^{\prime}-(S,Z)-Y$, with conditional distribution:
@@
This conditional distribution is always well-defined, given that $p_{S_{k}Z_{k}}$ has the same support as $p_{SZ}$. By assumption, $p_{S_{k}Z_{k}Y_{k}}\to p_{SZY}$; therefore,
@@
Conditional mutual information is (uniformly) continuous due to the (uniform) continuity of entropy [Theorem 17.3.3, 57]. Therefore,
$0=\lim_{k}[I(Q_{k};S_{k}|Y_{k})-I(Q_{k}^{\prime};S|Y)]\leq R-\lim_{k}I(Q_{k}^{\prime};S|Y),$ (46)
where we used Eq. (43). We also define another random variable $Q_{k}^{\alpha}$, which also obeys the Markov condition $Q_{k}^{\alpha}-(S,Z)-Y$, whose conditional distribution is defined in terms of the convex mixture:
@@
Here, $p_{U}(q)=1/|\mathcal{Q}|$ is a uniform distribution over an auxiliary independent random variable $U$ with outcomes $\mathcal{Q}$. From the convexity of conditional mutual information [58],
@@
Combining Eqs. (46), (47), and $R>0$ implies that $\alpha_{k}\to 1$, so
$p_{Q_{k}^{\alpha}SZY}-p_{Q_{k}^{\prime}SZY}\to 0.$
Combining with Eqs. (48), (45), and (43), along with continuity of conditional mutual information, gives
$I_{\text{RB}}(R)\geq\lim_{k}I(Q_{k}^{\alpha};Y|S)=\lim_{k}I(Q_{k}^{\prime};Y|S)\\ =\lim_{k}I(Q_{k};Y_{k}|S_{k})=\lim_{k}I_{\mathrm{\text{RB}}}^{k}(R).$
@@
00 0 $=\lim_{k}[I(Q;S|Y)-I(Q_{k}^{\prime\prime};S_{k}|Y_{k})]\leq R-\lim_{k}I(Q_{k}^{\prime\prime};S_{k}|Y_{k}).$ (51)
@@
Using the convexity of conditional mutual information, $I(U;S_{k};Y_{k})=0$, and the definition of $\alpha_{k}^{\prime}$, we have
@@
Combining Eqs. (51), (52), and $R>0$ implies $\alpha_{k}^{\prime}\to 1$; therefore,
$p_{Q_{k}^{\alpha\prime}S_{k}Z_{k}Y_{k}}-p_{Q_{k}^{\prime\prime}S_{k}Z_{k}Y_{k}}\to 0.$
Combining this with Eqs. (53), (50), and (42), along with the continuity of conditional mutual information, gives
$\lim_{k}I_{\mathrm{\text{RB}}}^{k}(R)\geq\lim_{k}I(Q_{k}^{\alpha\prime};Y_{k}|S_{k})\\ =\lim_{k}I(Q_{k}^{\prime\prime};Y_{k}|S_{k})=I(Q;Y|S)=I_{\text{RB}}(R).$